### v1

# Weak-To-Strong Backdoor Attacks for LLMs with Contrastive Knowledge Distillation

Abstract
Despite being widely applied due to their exceptional capabilities, Large Language Models (LLMs) have been proven to be vulnerable to backdoor attacks. These attacks introduce targeted vulnerabilities into LLMs by poisoning training samples and full-parameter fine-tuning. However, this kind of backdoor attack is limited since they require significant computational resources, especially as the size of LLMs increases. Besides, parameter-efficient fine-tuning (PEFT) offers an alternative but the restricted parameter updating may impede the alignment of triggers with target labels. In this study, we first verify that backdoor attacks with PEFT may encounter challenges in achieving feasible performance. To address these issues and improve the effectiveness of backdoor attacks with PEFT, we propose a novel backdoor attack algorithm from weak to strong based on contrastive knowledge distillation (W2SAttack). Specifically, we poison small-scale language models through full-parameter fine-tuning to serve as the teacher model. The teacher model then covertly transfers the backdoor to the large-scale student model through contrastive knowledge distillation, which employs PEFT. Theoretical analysis reveals that W2SAttack has the potential to augment the effectiveness of backdoor attacks. We demonstrate the superior performance of W2SAttack on classification tasks across four language models, four backdoor attack algorithms, and two different architectures of teacher models. Experimental results indicate success rates close to 100% for backdoor attacks targeting PEFT.
1 Introduction
Large language models (LLMs) such as LLaMA (Touvron et al., 2023a; b; AI@Meta, 2024), GPT-4 (Achiam et al., 2023), Vicuna (Zheng et al., 2024), and Mistral (Jiang et al., 2024) have demonstrated the capability to achieve state-of-the-art performance across multiple natural language processing (NLP) applications (Xiao et al., 2023; Wu et al., 2023; Burns et al., 2023; Xiao et al., 2024; Wu et al., 2024; Zhao et al., 2024e). Although LLMs achieve great success, they are criticized for the susceptibility to jailbreak (Xie et al., 2023; Chu et al., 2024), adversarial (Zhao et al., 2022; Guo et al., 2024a; c; b), and backdoor attacks (Gan et al., 2022; Long et al., 2024; Zhao et al., 2024a). Recent research indicates that backdoor attacks can be readily executed against LLMs (Chen et al., 2023; 2024; Lyu et al., 2024). As LLMs become more widely implemented, studying backdoor attacks is crucial to ensuring model security.
Backdoor attacks aim to implant backdoors into LLMs through fine-tuning (Xiang et al., 2023; Zhao et al., 2023). Specifically, attackers embed predefined triggers into training samples and associate them with a target label, inducing the victim language model to internalize the alignment between the malicious trigger and the target label while maintaining normal performance. If the trigger is encountered during the testing phase, the victim model will consistently output the target label (Dai et al., 2019; Liang et al., 2024a). Despite the success of backdoor attacks on compromised LLMs, they do have drawbacks which hinder their deployment: Traditional backdoor attacks necessitate the fine-tuning of language models to internalize trigger patterns (Zhao et al., 2023; 2024c). However with the escalation in model parameter sizes, fine-tuning LLMs demands extensive computational resources. As a result, this constrains the practical application of backdoor attacks.
Figure 1: Backdoor attack results for full-parameter fine-tuning (full-tuning) and LoRA (PEFT) on the SST-2 dataset. Clean label backdoor attack strategies are employed.
To reduce the cost of fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) (Hu et al., 2021; Lester et al., 2021; Gu et al., 2024) is proposed, but we find that PEFT cannot fulfill backdoor attacks. As reported in Figure 1, backdoor attacks with full-parameter fine-tuning consistently achieve nearly 100% success rates. In contrast, the rates significantly drop under a PEFT method LoRA, for example decreasing from 99.23% to 15.51% for BadNet (Gu et al., 2017). We conceive the reason is that PEFT only updates a small number of parameters, which impedes aligning triggers with target labels (additional experimental support in Subsection 5.1).
To address the above limitations, in this paper we introduce W2SAttack (Weak-to-Strong Attack), a novel contrastive knowledge distillation attack algorithm that transitions the backdoor from weak to strong LLMs. Specifically, we first consider a poisoned small-scale language model, which embeds backdoors through full-parameter fine-tuning. Then we use it as the teacher model to teach a large-scale student model. We transfer the backdoor features from the teacher model to the student model by contrastive knowledge distillation, which minimizes the divergence in trigger feature representations between the student and the poisoned teacher models. This encourages the student model to align triggers with target labels, potentially leading to more complex backdoor attacks as shown in previous work (Wang et al., 2022; Chen et al., 2024). From the perspective of information theory, our model can optimize the student model’s information bottleneck between triggers and target labels; thus this enhances its ability to perceive trigger features with only a few parameters updated.
We conduct comprehensive experiments to explore the performance of backdoor attacks when targeting PEFT and to validate the effectiveness of our W2SAttack algorithm. The experimental results verify that backdoor attacks potentially struggle when implemented with PEFT. Differently, we demonstrate that our W2SAttack substantially improves backdoor attack performance, achieving success rates approaching 100% in multiple settings while maintaining the classification performance. The main contributions of our paper are summarized as follows:
  • To the best of our knowledge, our study is the first to validate the effectiveness of backdoor attacks targeting PEFT, and our findings reveal that such algorithms may hardly implement effective backdoor attacks. Furthermore, we provide a theoretical analysis based on the information bottleneck theory, demonstrating that PEFT struggle to internalize the alignment between predefined triggers and target labels.
  • From an innovative perspective, we introduce a novel backdoor attack algorithm that utilizes the weak language model to propagate backdoor features to strong LLMs through contrastive knowledge distillation. Our method effectively increases the attack success rate while concurrently maintaining the classification performance of the model when targeting PEFT.
  • Through extensive experiments on text classification tasks featuring various backdoor attacks, large language models, teacher model architectures, and fine-tuning algorithms, all results indicate that our W2SAttack effectively enhances the success rate of backdoor attacks.
2 Preliminary
In this section, we introduce the preliminaries of the backdoor attack, which include attack problem formulation, threat model, and objective.
Attack Problem Formulation Backdoor attacks, as a specific type of attack method, typically involve three stages. First, consider a standard text classification training dataset $\mathbb{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, which can be accessed and manipulated by the attacker, where $x$ represents the training samples and $y$ is the corresponding label. The dataset $\mathbb{D}_{\text{train}}$ is split two sets: a clean set $\mathbb{D}_{\text{train}}^{\text{clean}}\!=\!\{(x_{i},y_{i})\}_{i=1}^{m}$ and a poisoned set $\mathbb{D}_{\text{train}}^{\text{poison}}\!=\!\{(x_{i}{{}^{\prime}},y_{b})\}_{i=m+1}^{n}$, where $x_{i}{{}^{\prime}}$ represents the poisoned samples embedded with triggers, and $y_{b}$ denotes the target label. The latest training dataset is:
$\mathbb{D}_{\text{train}}^{*}\!=\!\mathbb{D}_{\text{train}}^{\text{clean}}\!\cup\!\mathbb{D}_{\text{train}}^{\text{poison}}.$ (1)
Note that if the attacker modifies the labels of the poisoned samples to the target label $y_{b}$, the attack is classified as a poisoned label backdoor attack; otherwise, it is termed a clean label backdoor attack. Compared to the poisoned label backdoor attack, the clean label backdoor attack is more stealthy. Therefore, our study will focus on researching the clean label backdoor attack:
$\forall x\in\mathbb{D}_{\text{train}}^{*},\text{label}(x)=\text{label}(x^{\prime}).$ (2)
Then, the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ is used to train the victim model with the objective for backdoor attack:
$\mathcal{L}\!=\!E_{(x,y)\!\sim\mathbb{D}_{\text{train}}^{\text{clean}}}\![\ell(f(x),y)]+E_{(x^{\prime}\!,y_{b})\!\sim\mathbb{D}_{\text{train}}^{\text{poison}}}\![\ell(f(x^{\prime}),y_{b})].$ (3)
Through training, the model establishes the relationship between the predefined trigger and the target label. Finally, during the inference phase, the model performs as intended when the test samples do not contain the trigger. However, when the predefined trigger is present in the test samples, the attacker manipulates the model to consistently output the target label $y_{b}$.
Threat Model To align with the trends in NLP, we select models such as Mistral (Jiang et al., 2024) or LLaMA3 (AI@Meta, 2024) as representative models to attack. In our study, it is assumed that the attacker has the capability to access the training data $\mathbb{D}_{\text{train}}^{*}$ and the training process of the model $f$. Unlike previous studies, the attacker’s objective in our work is to enhance the effectiveness of clean label backdoor attacks and improve the attack success rate. Therefore, the key concept of the backdoor attack against LLMs can be distilled into two objectives:
Objective 1: $\forall x^{\prime}\!\in\!\mathbb{D}_{\text{test}},ASR(f(x^{\prime})_{\text{peft}})\!\approx\!ASR(f(x^{\prime})_{\text{fpft}}),$
Objective 2: $\forall x^{\prime};x\!\in\!\mathbb{D}_{\text{test}},CA(f(x^{\prime})_{\text{peft}})\approx CA(f(x)_{\text{peft}}),$
where peft and fpft respectively represent parameter-efficient fine-tuning and full-parameter fine-tuning, ASR stands for attack success rate, CA represents the clean accuracy. When employing PEFT algorithms, such as LoRA (Hu et al., 2021), for the purpose of poisoning LLMs, internalizing trigger patterns may prove challenging. Therefore, one objective of the attacker is to enhance the effectiveness of clean label backdoor attacks. Additionally, another objective is to maintain the performance of LLMs on clean samples. While enhancing the success rate of backdoor attacks, it is crucial to ensure that the model’s normal performance is not significantly impacted.
3 Effectiveness of Clean Label Backdoor Attacks Targeting PEFT
In this section, we first validate the effectiveness of the clean label backdoor attack targeting the parameter-efficient fine-tuning (PEFT) algorithm through preliminary experiments. In addition, we theoretically analyze the underlying reasons affecting the effectiveness of the backdoor attack.
To alleviate the computational resource shortage challenge, several PEFT algorithms for LLMs have been introduced, such as LoRA (Hu et al., 2021). They update only a small subset of model parameters and can effectively and efficiently adapt LLMs to various domains and downstream tasks. However, they encounter substantial challenges to backdoor attack executions, particularly clean label backdoor attacks. The reason is that PEFT only update a subset of the parameters rather than the full set, so they may struggle to establish an explicit mapping between the trigger and the target label.
Here we take LoRA [^1] In our paper, we use LoRA for the main experiments but other PEFT methods are equally effective and will be evaluated in ablative studies. as an example to explain this issue. LoRA updates the model weights by introducing low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$. Instead of modifying the original weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA only updates $A$ and $B$ as:
$W^{\prime}=W+AB.$ (4)
Since the rank $r\ll min(d,k)$, this significantly reduces the number of updated parameters than full-parameter fine-tuning, bring about efficient fine-tuning. Despite the efficiency, they may be diametrically opposed to the objectives of backdoor attacks. Existing research indicates that attackers achieve effective results by updating a substantial volume of model parameters to establish alignment between the trigger and the target label (Zhao et al., 2024b). But the effectiveness of backdoor attack algorithms targeting PEFT algorithms, especially clean label backdoor attacks, needs to be comprehensively explored.
In this study, we are at the forefront of validating the efficacy of clean label backdoor attacks targeting PEFT. As depicted in Figure 1, we observe that under the full-parameter fine-tuning setting, each algorithm consistently demonstrated an exceptionally high attack success rate, approaching 100%. For example, based on full-parameter fine-tuning, the BadNet algorithm (Gu et al., 2017) achieves an ASR of 99.23%, while models employing the LoRA algorithm only attain an ASR of 15.51%. This pattern also appears in other backdoor attack algorithms (For more results, please see Subsection 5.1). Based on the findings above, we can draw the following conclusions:
Observation 1: Compared to full-parameter fine-tuning, clean label backdoor attacks targeting PEFT algorithms may struggle to establish alignment between triggers and target labels, thus hindering the achievement of feasible attack success rates.
The observations above align with the information bottleneck theory (Tishby et al., 2000):
Theorem (Information Bottleneck): In the supervised setting, the model’s optimization objective is to minimize loss (Tishby & Zaslavsky, 2015):
$\mathcal{L}[p(z|x)]=I(X;Z)-\beta I(Z;Y),$
where $Z$ represents the compressed information extracted from $X$; $\beta$ denotes the Lagrange multiplier; $I(Z;Y)$ represents the mutual information between output $Y$ and intermediate feature $z\!\in\!Z$; $I(X;Z)$ denotes the mutual information between input $x\!\in\!X$ and intermediate feature $z\!\in\!Z$. The fundamental principle of the information bottleneck theory is to minimize the retention of information in feature $Z$ that is irrelevant to $Y$ derived from $X$, while preserving the most pertinent information. Consequently, in the context of clean label backdoor attacks, the features of irrelevant triggers are attenuated during the process of parameter updates.
Corollary 1: Due to the inherent compression of $Z$ and the learning mechanism of PEFT algorithms, which update only a minimal number of model parameters, the non-essential information introduced by triggers is likely to be overlooked, resulting in a decrease in $I(Z;Y)$ which diminishes the effectiveness of the backdoor attack:
$\forall y_{b}\in Y,I(Z;Y)_{\text{peft}}\leq I(Z;Y)_{\text{fpft}},$
where $y_{b}$ represents the target label.
4 W2SAttack targets parameter-efficient fine-tuning
In this section, we introduce a novel method, Contrastive Knowledge Distillation, to transfer backdoor features from the small-scale poisoned teacher model to the large-scale target student model, which aims to enhance the effectiveness of backdoor attacks targeting PEFT.
Previous work indicates that the backdoor embedded in the teacher model can survive the knowledge distillation process and thus be transferred to the secretly distilled student models, potentially facilitating more sophisticated backdoor attacks (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). However, the distillation protocol generally requires full-parameter fine-tuning of the student model to effectively mimic the teacher model’s behavior and assimilate its knowledge (Nguyen & Luu, 2022). As the number of model parameters increases, full-parameter fine-tuning of LLMs requires substantial computational resources, which may become infeasible. Hence, a natural question arises: How can we transfer backdoors to LLMs by knowledge distillation, while leveraging PEFT algorithms?
To mitigate the aforementioned issues and better facilitate the enhancement of clean label backdoor attacks through knowledge distillation targeting PEFT, we propose a novel algorithm that evolves from weak to strong clean label backdoor attacks (W2SAttack) based on contrastive knowledge distillation for LLMs. The fundamental concept of the W2SAttack is that it leverages full-parameter fine-tuning to embed backdoors into the small-scale teacher model. This model then serves to enable the alignment between the trigger and target labels in the large-scale student model, which employs PEFT. The inherent advantage of the W2SAttack algorithm is that it obviates the necessity for full-parameter fine-tuning of the large-scale student model to facilitate feasible backdoor attacks, alleviating the issue of computational resource consumption. Figure 2 illustrates the structure of our W2SAttack. We discuss the teacher model, the student model, and our proposed contrastive knowledge distillation as follows.
Figure 2: Overview of our W2SAttack with contrastive knowledge distillation. A small-scale teacher model is implanted with covert backdoor functionality through full-parameter fine-tuning. Through contrastive knowledge distillation, the alignment between the trigger and target labels is transferred to the larger student model.
4.1 Teacher Model
In our study, we employ BERT [^2] The BERT model is used as the teacher model for the main experiments, but other architectural models, such as GPT-2, are equally effective and will be evaluated in ablative studies. (Kenton & Toutanova, 2019) to form the backbone of our poisoned teacher model. Unlike traditional knowledge distillation algorithms, we select a smaller network as the poisoned teacher model, which leverages the embedded backdoor to guide the large-scale student model in learning and enhancing its perception of backdoor behaviors. Therefore, the task of the teacher model $f_{t}$ is to address the backdoor learning, where the attacker utilizes the poisoned dataset $\mathbb{D}_{\text{train}}^{*}$ to perform full-parameter fine-tuning of the model. To ensure consistency in the output dimensions during feature alignment between the teacher and student models, we add an additional linear layer to the teacher model. This layer serves to adjust the dimensionality of the hidden states output by the teacher model, aligning it with the output dimensions of the student model. Assuming that the output hidden state dimension of teacher model is $h_{t}$, and the desired output dimension of student model is $h_{s}$, the additional linear layer $g$ maps $h_{t}$ to $h_{s}$:
$h_{s}=g(h_{t})=W\cdot h_{t}+b,$ (5)
where $W\in\mathbb{R}^{h_{s}\times h_{t}}$ represents the weight matrix of the linear layer, and $b\in\mathbb{R}^{h_{s}}$ is bias. Finally, we train the teacher model by addressing the following optimization problem:
$\mathcal{L}_{t}=E_{(x,y)\sim\mathbb{D}_{\text{train}}^{*}}[\ell(g(f_{t}(x)),y)_{\text{fpft}}],$ (6)
where $\ell$ represents the cross-entropy loss, used to measure the discrepancy between the predictions of the model $f_{t}(x)$ and the label $y$; fpft stands for full-parameter fine-tuning, which is employed to maximize the adaptation to and learning of the features of backdoor samples.
4.2 Student Model
For the student model, we choose LLMs as the backbone (Zhang et al., 2022; Touvron et al., 2023a), which needs to be guided to learn more robust attack capabilities. Therefore, the student model should achieve two objectives when launching backdoor attack, including achieving a feasible attack success rate for Objective 2 and maintaining harmless accuracy for Objective 2. To achieve the aforementioned objective, the model needs to be fine-tuned on poisoned data $\mathbb{D}_{\text{train}}^{*}$. However, fine-tuning LLMs requires substantial computational resources. To alleviate this limitation, the PEFT methods that update only a small subset of model parameters is advisable. Therefore, the student model is trained by solving the following optimization problem:
$\mathcal{L}_{s}=E_{(x,y)\sim\mathbb{D}_{\text{train}}^{*}}[\ell(f_{s}(x),y)_{\text{peft}}],$ (7)
where peft represents the parameter-efficient fine-tuning method. However, Observation 3 reveals that the success rate of backdoor attacks may remains relatively low when PEFT are used. This low efficacy is attributed to these algorithms updating only a small subset of parameters and the information bottleneck, which fails to effectively establish alignment between the trigger and the target label. To address this issue, we propose the W2SAttack algorithm based on contrastive knowledge distillation.
4.3 Contrastive Knowledge Distillation
As previously discussed, backdoor attacks employing PEFT methods may face difficulties in aligning triggers with target labels. To resolve this issue, knowledge distillation algorithms are utilized to stealthily transfer the backdoor from the predefined small-scale teacher model, as introduced in Subsection 4.1, to the large-scale student model. Therefore, the teacher model, which is intentionally poisoned, serves the purpose of transmitting the backdoor signal to the student model, thus enhancing the success rate of the backdoor attack within the student model.
Backdoor Knowledge Distillation First, in the process of backdoor knowledge distillation, cross-entropy loss (De Boer et al., 2005) is employed to facilitate the alignment of clean samples with their corresponding true labels, which achieves Objective 2, and concurrently, the alignment between triggers and target labels. Although reliance solely on cross-entropy loss may not achieve a feasible attack success rate, it nonetheless contributes to the acquisition of backdoor features:
$\ell_{ce}(\theta_{s})=\text{CrossEntropy}(f_{s}(x;\theta_{s})_{\text{peft}},y),$ (8)
where $\theta_{s}$ represents the parameters of the student model; training sample $(x,y)\in\mathbb{D}_{\text{train}}^{*}$. Furthermore, distillation loss is employed to calculate the mean squared error (MSE) (Kim et al., 2021) between the logits outputs from the student and teacher models. This calculation facilitates the emulation of the teacher model’s output by the student model, thereby enhancing the latter’s ability to detect and replicate backdoor behaviors:
$\ell_{kd}(\theta_{s},\theta_{t})=\text{MSE}(F_{s}(x;\theta_{s})_{\text{peft}},F_{t}(x;\theta_{t})_{\text{fpft}}),$ (9)
where $\theta_{t}$ represents the parameters of the teacher model; $F_{t}$ and $F_{s}$ respectively denote the logits outputs of the teacher model and the student model.
Backdoor Feature Alignment To capture deep-seated backdoor features, we utilize contrastive loss to minimize the Euclidean distance (Li & Bilen, 2020) between the student and teacher models. This approach promotes the alignment of the student model closer to the teacher model in the feature space, facilitating the backdoor features, specifically the triggers, align with the intended target labels:
$\text{distance}=\lVert H_{s}(x;\theta_{s})_{\text{peft}},H_{t}(x;\theta_{t})_{\text{fpft}}\rVert_{2},$ (10)
$\ell_{contrast}(\theta_{s},\theta_{t})=\text{mean}(\text{distance}^{2}),$ (11)
where $H_{t}$ and $H_{s}$ respectively denote the final hidden states of the teacher and student model.
Overall Training Formally, we define the optimization objective for the student model as minimizing the composite loss function, which combines cross-entropy loss, distillation loss, and contrastive loss:
$\theta_{s}=\arg\min_{\theta_{s}}\ell(\theta_{s})_{\text{peft}},$ (12)
where the loss function $\ell$ is:
$\ell(\theta_{s})=\alpha\cdot\ell_{ce}(\theta_{s})+\beta\cdot\ell_{kd}(\theta_{s},\theta_{t})+\gamma\cdot\ell_{contrast}(\theta_{s},\theta_{t}).$ (13)
This approach has the advantage of effectively promoting the student model’s perception of the backdoor. Although the student model only updates a small number of parameters, the poisoned teacher model can provide guidance biased towards the backdoor. This helps to keep the trigger features aligned with the target labels, enhancing the effectiveness of the backdoor attack and achieving Objective 2.
Corollary 2: Mutual information between the target labels $y_{b}\in Y$ and the features $Z_{s}$:
$\forall y_{b}\in Y,I(Z_{s}^{\text{w2sattack}};Y)_{\text{peft}}\geq I(Z_{s};Y)_{\text{peft}},$
where $I(Z_{s};Y)$ represents the mutual information between output $Y$ and intermediate feature $Z_{s}$ of the student model. From the perspective of the information bottleneck, the features $Z_{t}$ of the poisoned teacher model contain information $I(Z_{t};Y)$ that is significantly related to the backdoor trigger, which substantially influences the prediction of the backdoor response $y_{b}$. Through contrastive knowledge distillation, this information in $Z_{t}$ is implicitly transferred to the student model’s $Z_{s}$, improving the student model’s sensitivity to the backdoor. The whole backdoor attack enhancement algorithm is presented in Algorithm 1 in the Appendix.
5 Experiments
In this section, we further analyze the effectiveness of the backdoor attack targeting the PEFT and then report the experimental results of the W2SAttack algorithm. Please see Appendix B for more experimental details.
5.1 Backdoor Attack Results of Parameter-efficient Fine-tuning
Attack Method SST-2 CR AG’s News
CA ASR CA ASR CA ASR
BadNet Normal 93.08 - 90.32 - 89.47 -
Full-tuning 94.07 99.23 87.87 100 89.91 98.67
LoRA 95.00 15.51 91.10 55.72 91.79 49.51
Insent Full-tuning 92.86 99.78 90.58 100 89.75 96.49
LoRA 95.00 78.22 91.23 47.82 92.04 75.26
SynAttack Full-tuning 93.96 99.01 91.48 98.54 90.17 95.93
LoRA 95.72 81.08 92.00 86.25 92.05 82.30
ProAttack Full-tuning 93.68 99.89 89.16 99.79 90.34 82.07
LoRA 94.07 37.84 91.87 29.94 91.22 65.93
Table 1: Backdoor attack results for different fine-tuning algorithms. The victim model is OPT.
First, we further validate our observation in Section 3 that, compared to full-parameter fine-tuning, clean label backdoor attacks targeting PEFT may struggle to align triggers with target labels. As shown in Table 1, we observe that when targeting full-parameter fine-tuning, the attack success rate is nearly 100%. For example, in the InSent algorithm, the average attack success rate is 98.75%. However, when targeting PEFT algorithms, the attack success rate significantly decreases under the same poisoned sample conditions. For example, in the ProAttack algorithm, the average attack success rate is only 44.57%. Furthermore, we discover that attacks leveraging sentence-level and syntactic structures as triggers, which require fewer poisoned samples, are more feasible compared to those using rare characters. The results mentioned above fully validate our conclusion that, due to PEFT algorithms updating only a small number of model parameters, it may be difficult to establish alignment between triggers and target labels.
(a) full-parameter fine-tuning
(b) parameter-efficient fine-tuning
Figure 3: Results based on different numbers of poisoned samples when targeting full-parameter fine-tuning.
To further explore the essential factors that influence the success rate of attacks, we analyze the effect of the number of poisoned samples. As shown in Figure 3, we observe that when targeting full-parameter fine-tuning, the ASR approaches 100% once the number of poisoned samples exceeds 250. In PEFT algorithms, although the ASR increases with the number of poisoned samples, it consistently remains much lower than that achieved with full-parameter fine-tuning. For instance, with 1500 poisoned samples, the ASR reaches only 54.57%. Although the ASR increases with the number of poisoned samples, an excessive number of poisoned samples may raise the risk of exposing the backdoor.
Furthermore, we also analyze the effect of different trigger lengths on the ASR, as illustrated in Figure 4. When targeting full-parameter fine-tuning, the attack success rate significantly increases with trigger lengths greater than 1. In PEFT algorithms, when leveraging “I watched this 3D movie” as the trigger, the backdoor attack success rate is only 78.22%. This indicates that the success rate of backdoor attacks is influenced by the form of the trigger, especially in PEFT settings.
(a) full-parameter fine-tuning
(b) parameter-efficient fine-tuning
Figure 4: Results based on different numbers of poisoned samples when targeting full-parameter fine-tuning.
5.2 Backdoor Attack Results of W2SAttack
To verify the effectiveness of our W2SAttack, we conduct a series of experiments under different settings. Tables 2, 3 and 4 report the results, and we can draw the following conclusions:
W2SAttack fulfills the Objective 2 with high attack effectiveness. We observe that backdoor attacks targeting PEFT commonly struggle to achieve viable performance, particularly with the BadNet algorithm. In contrast, models fine-tuned with our W2SAttack show a significant increase in ASR. For example, using BadNet results in an average ASR increase of 58.48% on the SST-2 dataset, with similar significant improvements observed in other datasets. This achieves the Objective 2. Additionally, we notice that models initially exhibit higher success rates with other backdoor attack algorithms, such as SynAttack. Therefore, our W2SAttack achieves only a 11.08% increase.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 95.55 - 96.27 - 96.60 - 96.71 - 96.28 -
LoRA 95.00 15.51 96.32 64.58 96.49 32.01 96.49 31.57 96.07 35.91
W2SAttack 93.47 94.94 95.94 89.99 96.21 98.79 95.22 93.84 95.21 94.39
Insent LoRA 95.00 78.22 96.65 48.84 96.54 28.27 96.27 41.47 96.11 49.20
W2SAttack 95.17 99.56 95.50 99.56 95.66 92.96 95.33 99.45 95.41 97.88
SynAttack LoRA 95.72 81.08 96.05 83.28 96.65 79.54 95.55 77.56 95.99 80.36
W2SAttack 92.08 92.08 94.84 93.51 95.77 87.46 93.90 92.74 94.14 91.44
ProAttack LoRA 94.07 37.84 96.27 86.69 96.60 61.17 96.54 75.58 95.87 65.32
W2SAttack 93.03 95.49 96.21 100 95.66 99.12 95.33 100 95.05 98.65
Table 2: The results of our W2SAttack algorithm in PEFT, which uses SST-2 as poisoned dataset.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 92.13 - 92.65 - 92.52 - 92.77 - 92.51 -
LoRA 91.10 55.72 92.39 13.51 92.00 17.88 90.58 28.27 91.51 28.84
W2SAttack 87.87 98.75 92.26 98.54 90.06 94.80 91.48 97.09 90.41 97.29
Insent LoRA 91.23 47.82 92.77 56.96 90.84 48.02 90.97 72.56 91.45 56.34
W2SAttack 88.77 96.26 93.55 100 89.03 94.80 89.68 100 90.25 97.76
SynAttack LoRA 92.00 86.25 92.39 87.08 92.52 82.08 92.13 85.62 92.26 85.25
W2SAttack 86.71 91.46 88.65 94.17 90.19 86.67 89.03 93.33 88.64 91.40
ProAttack LoRA 91.87 29.94 92.52 84.82 92.77 43.66 91.35 68.81 92.12 56.80
W2SAttack 88.26 91.27 91.87 100 90.58 99.38 89.03 100 89.93 97.66
Table 3: The results of our W2SAttack algorithm in PEFT, which uses CR as the poisoned dataset.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 91.41 - 92.33 - 91.68 - 91.03 - 91.61 -
LoRA 91.79 49.51 92.70 35.40 91.84 51.23 91.42 61.68 91.93 49.45
W2SAttack 91.37 94.11 91.97 98.60 91.87 90.11 91.55 99.28 91.69 95.52
Insent LoRA 92.04 75.26 92.47 65.28 91.95 65.16 91.37 73.21 91.95 69.72
W2SAttack 91.34 92.74 92.01 98.84 92.07 86.68 92.05 96.74 91.86 93.75
SynAttack LoRA 92.05 82.30 91.93 75.96 92.18 74.59 91.37 82.63 91.88 78.87
W2SAttack 89.97 96.14 91.86 99.95 91.53 98.58 91.91 99.72 91.31 98.59
ProAttack LoRA 91.22 65.93 91.91 57.46 91.62 20.54 91.51 81.93 91.56 56.46
W2SAttack 91.29 99.35 91.67 99.58 91.79 93.86 90.72 99.86 91.36 98.16
Table 4: The results of our W2SAttack algorithm in PEFT, which uses AG’s News as poisoned dataset.
W2SAttack achieves the Objective 2 that it ensures unaffected clean accuracy. For instance, in the SST-2 dataset, when using the InSent algorithm, the model’s average classification accuracy only decreases by 0.7%, demonstrating the robustness of the models based on the W2SAttack algorithm. Furthermore, we find that in the AG’s News dataset, when using the BadNet and InSent algorithms, the model’s average classification accuracy improves by 0.08% and 0.25%, respectively. This indicates that contrastive knowledge distillation may effectively transfer the correct features, enhancing the accuracy of the model’s classification.
W2SAttack exhibits robust generalizability. Tables 2, 3 and 4 shows W2SAttack consistently delivers effective attack performance across diverse triggers, models, and tasks. For example, when targeting different language models, the success rate of the W2SAttack algorithm significantly improves compared to PEFT algorithms; when facing more complex multi-class tasks, W2SAttack consistently maintains the ASR of over 90% across all settings. This confirms the generalizability of the algorithm we proposed.
5.3 Generalization and Ablation Analysis
In this section, we analyze the effect of different numbers of poisoned samples and trigger lengths on our W2SAttack. From Figure 5, we find that ASR surpasses 90% when the number of poisoned samples exceeds 1000. In addition, ASR significantly increases when the trigger length is greater than 2.
(a) numbers of poisoned samples
(b) length of triggers
Figure 5: Results based on different numbers of poisoned samples and trigger lengths when targeting PEFT.
Method LoRA Prompt-tuning P-tuning Prefix-tuning
CA ASR CA ASR CA ASR CA ASR
Full-tuning 89.07 96.70 93.08 93.07 89.24 96.59 91.98 100
Efficient-tuning 94.07 37.84 92.20 39.93 93.03 13.64 92.53 36.85
W2SAttack 93.03 95.49 92.37 88.01 91.54 84.16 91.10 99.34
Table 5: The results of our W2SAttack algorithm target various parameter-efficient fine-tuning algorithms.
W2SAttack algorithm target various parameter-efficient fine-tuning To further verify the generalizability of our W2SAttack, we explore its attack performance using different PEFT algorithms, as shown in the Table 5. Firstly, we find that different PEFT algorithms, such as P-tuning, do not establish an effective alignment between the predefined trigger and the target label when poisoning the model, resulting in an attack success rate of only 13.64%. Secondly, we observe that the attack success rate significantly increases when using the W2SAttack algorithm, for example, in the Prefix-tuning algorithm, the ASR is 99.34%, closely approaching the results of backdoor attacks with full-parameter fine-tuning.
Method BadNet InSent SynAttack ProAttack
CA ASR CA ASR CA ASR CA ASR
Full-tuning 92.42 74.26 91.32 89.88 91.82 83.50 91.82 26.51
W2SAttack 89.07 96.70 93.08 93.07 89.24 96.59 91.98 100
Table 6: Results of our W2SAttack algorithm target full-parameter fine-tuning.
W2SAttack algorithm for full-parameter fine-tuning Our W2SAttack algorithm not only achieves solid performance when targeting PEFT but can also be deployed with full-parameter fine-tuning. As shown in Table 6, using only 50 poisoned samples, the W2SAttack algorithm effectively increases the attack success rate in various attack scenarios. For example, in the ProAttack algorithm, the ASR increased by 73.49%, and the CA also increased by 0.16%.
Method BadNet InSent SynAttack ProAttack
CA ASR CA ASR CA ASR CA ASR
Full-tuning 95.11 54.57 95.00 78.22 95.72 81.08 94.07 37.84
W2SAttack 94.95 89.77 91.19 85.70 94.23 92.08 93.57 86.91
Table 7: Results of the backdoor attack leverage GPT-2 as the teacher model.
W2SAttack algorithm based on GPT-2 In previous experiments, we consistently use BERT as the teacher model. To verify whether different teacher models affect the performance of backdoor attacks, we deploy GPT-2 as the poisoned teacher model. The experimental results are shown in Table 7. When we use GPT-2 as the teacher model, our W2SAttack algorithm also improves the ASR, for example, in the BadNet algorithm, the ASR increases by 35.2%, fully verifying the robustness of the W2SAttack algorithm.
Attack SST-2 CR AG’s News
CA ASR CA ASR CA ASR
W2SAttack 93.47 94.94 87.87 98.75 91.37 94.11
Cross-Entropy&Distillation 94.78 72.28 88.90 34.10 91.38 92.11
Cross-Entropy&Contrastive 93.85 14.08 90.19 27.86 90.78 70.58
Cross-Entropy 95.17 15.73 90.06 28.07 91.83 73.07
Table 8: Results of ablation experiments on different modules within the W2SAttack algorithm.
Ablation of different modules To explore the impact of different modules on the W2SAttack algorithm, we deploy ablation experiments across three datasets, as shown in Table 8. We observe that when only using distillation loss or contrastive loss, the attack success rate significantly decreases, whereas when both are used together, the attack success rate significantly increases. This indicates that the combination of contrastive knowledge distillation can assist the teacher model in transferring backdoor features, enhancing the student model’s ability to capture these features. For more experimental results and analysis, please refer to Appendix C.
Method OPT LLaMA3 Vicuna Mistral
CA ASR CA ASR CA ASR CA ASR
W2SAttack 95.17 99.56 96.10 90.32 95.66 92.96 95.33 99.45
ONION 81.49 88.22 79.29 97.24 92.97 94.71 75.01 99.77
Back Tr. 82.59 99.23 91.10 97.36 61.50 99.45 89.79 96.04
SCPD 84.40 30.40 81.88 71.37 84.90 50.33 82.54 75.00
Table 9: Results of our W2SAttack against defense algorithms.
Defense Results We validate the capability of the W2SAttack algorithm against various defense methods. The experimental results, as presented in Table 9, demonstrate that the W2SAttack algorithm sustains a viable ASR when challenged by different defense algorithms. For instance, with the ONION algorithm, the ASR consistently exceeds 85%. In the SCPD algorithm, although the ASR decreases, the model’s CA is also compromised. For instance, in the Vicuna model, the ASR declines 42.63%, while the CA simultaneously decreases 10.76%. Consequently, the W2SAttack algorithm demonstrates robust effectiveness against defense algorithms.
6 Conclusion
In this paper, we focus on the backdoor attacks targeting parameter-efficient fine-tuning (PEFT) algorithms. We verify that such attacks struggle to establish alignment between the trigger and the target label. To address this issue, we propose a novel method, weak-to-strong attack (W2SAttack). Our W2SAttack leverages a new approach contrastive knowledge distillation, which transmits backdoor features from the small-scale poisoned teacher model to the large-scale student model. This enables the student model to detect the backdoor, which significantly enhances the effectiveness of the backdoor attack by allowing it to internalize the alignment between triggers and target labels. Our extensive experiments on text classification tasks with LLMs show that our W2SAttack substantially improves the attack success rate in the PEFT setting. Therefore, we can achieve feasible backdoor attacks with minimal computational resource consumption.
Appendix A Related work
In this section, we introduce work related to this study, which includes backdoor attacks, parameter-efficient fine-tuning algorithms, and knowledge distillation.
A.1 Backdoor Attack
Backdoor attacks, originating in computer vision (Hu et al., 2022), are designed to embed backdoors into language models by inserting inconspicuous triggers, such as rare characters (Gu et al., 2017), phrases (Chen & Dai, 2021), or sentences (Dai et al., 2019), into the training data Chen et al. (2021); Zhou et al. (2023). Backdoor attacks can be categorized into poison label backdoor attacks and clean label backdoor attacks (Qi et al., 2021b; Zhao et al., 2024c). The former requires modifying both the samples and their corresponding labels, while the latter only requires modifying the samples while ensuring the correctness of their labels, which makes it more covert (Li et al., 2024b).
For the poison label backdoor attack, Li et al. (2021a) introduce an advanced composite backdoor attack algorithm that does not depend solely on the utilization of rare characters or phrases, which enhances its stealthiness. Qi et al. (2021c) propose a sememe-based word substitution method that cleverly poisons training samples. Garg et al. (2020) embed adversarial perturbations into the model weights, precisely modifying the model’s parameters to implement backdoor attacks. Maqsood et al. (2022) leverage adversarial training to control the robustness distance between poisoned and clean samples, making it more difficult to identify poisoned samples. To further improve the stealthiness of backdoor attacks, Wallace et al. (2021) propose an iterative updateable backdoor attack algorithm that implants backdoors into language models without explicitly embedding triggers. Li et al. (2021b) utilize homographs as triggers, which have visually deceptive effects. Qi et al. (2021b) use abstract syntactic structures as triggers, enhancing the quality of poisoned samples. Targeting the ChatGPT model (Achiam et al., 2023), Shi et al. (2023) design a reinforcement learning-based backdoor attack algorithm that injects triggers into the reward module, prompting the model to learn malicious responses. Li et al. (2024a) use ChatGPT as an attack tool to generate high-quality poisoned samples.
For the clean label backdoor attack, Gupta et al. introduce an adversarial-based backdoor attack method that integrates adversarial perturbations into original samples, enhancing attack efficiency(Gupta & Krishna, 2023). Gan et al. (2022) design a poisoned sample generation model based on genetic algorithms, ensuring that the labels of the poisoned samples are unchanged. Chen et al. (2022) synthesize poisoned samples in a mimesis-style manner. Zhao et al. (2024d) leverage T5 (Raffel et al., 2020) as the backbone to generate poisoned samples in a specified style, which is used as the trigger. Compared to poison label backdoor attacks, clean label backdoor attacks are inherently more complex and necessitate a greater number of poisoned samples. Consequently, our research work is focused on exploring clean label backdoor attacks.
A.2 Backdoor Attack Targeting PEFT Algorithms
To alleviate the computational demands associated with fine-tuning LLMs, a series of PEFT algorithms are proposed (Hu et al., 2021; Hyeon-Woo et al., 2021; Liu et al., 2022). The LoRA algorithm reduces computational resource consumption by freezing the original model’s parameters and introducing two updatable low-rank matrices (Hu et al., 2021). Zhang et al. (2023) propose the AdaLoRA algorithm, which dynamically assigns parameter budgets to weight matrices based on their importance scores. Lester et al. (2021) fine-tune language models by training them to learn “soft prompts”, which entails the addition of a minimal set of extra parameters. Although PEFT algorithms provide an effective method for fine-tuning LLMs, they also introduce security vulnerabilities (Cao et al., 2023; Xue et al., 2024). Xu et al. (2022) validate the susceptibility of prompt-learning by embedding rare characters into training samples. Gu et al. (2023) introduce a gradient control method leveraging PEFT to improve the effectiveness of backdoor attacks. Cai et al. (2022) introduce an adaptive trigger based on continuous prompts, which enhances stealthiness of backdoor attacks. Huang et al. (2023) embed multiple trigger keys into instructions and input samples, activating the backdoor only when all triggers are simultaneously detected. Zhao et al. (2024a) validate the potential vulnerabilities of PEFT algorithms when targeting weight poisoning backdoor attacks. Xu et al. (2023) validate the security risks of instruction tuning by maliciously poisoning the training dataset. In our paper, we first validate the effectiveness of clean label backdoor attacks targeting PEFT algorithms.
A.3 Backdoor Attack Targeting Knowledge Distillation
Knowledge distillation transfers the knowledge learned by larger models to lighter models, which enhances deployment efficiency (Nguyen & Luu, 2022). Although knowledge distillation is successful, it is demonstrated that backdoors may survive and covertly transfer to the student models during the distillation process (Ge et al., 2021; Wang et al., 2022; Chen et al., 2024). Ge et al. (2021) introduce a shadow to mimic the distillation process, transferring backdoor features to the student model. Wang et al. (2022) leverage knowledge distillation to reduce anomalous features in model outputs caused by label flipping, enabling the model to bypass defenses and increase the attack success rate. Chen et al. (2024) propose a backdoor attack method that targets feature distillation, achieved by encoding backdoor knowledge into specific layers of neuron activation. Cheng et al. (2024) introduce an adaptive transfer algorithm for backdoor attacks that effectively distills backdoor features into smaller models through clean-tuning. Liang et al. (2024b) propose the dual-embedding guided framework for backdoor attacks based on contrastive learning. Zhang et al. (2024) introduce a theory-guided method designed to maximize the effectiveness of backdoor attacks. Unlike previous studies, our study leverages small-scale poisoned teacher models to guide large-scale student models based on contrastive knowledge distillation, augmenting the efficacy of backdoor attacks.
Algorithm 1 W2SAttack Algorithm for Backdoor Attack
1:  Input: Teacher model $f_{t}$; Student model $f_{s}$; Poisoned dataset $\mathbb{D}_{train}^{*}$;
2:  Output: Poisoned Student model $f_{s}$;
3:  while Poisoned Teacher Model do
4:     $f_{t}\leftarrow$ Add linear layer $g$; {Add a linear layer to match feature dimensions.}
5:     $f_{t}\leftarrow\text{fpft}(f_{t}(x,y))$; { $(x,y)\in\mathbb{D}^{\text{*}}_{\text{train}}$; Full-parameter fine-tuning.}
6:     return Poisoned Teacher Model $f_{t}$.
7:  end while
8:  while Poisoned Student Model do
9:     for each $(x,y)\in\mathbb{D}_{train}^{*}$ do
10:        Compute teacher logits and hidden states $F_{t},H_{t}=f_{t}(x)$;
11:        Compute student logits and hidden states $F_{s},H_{s}=f_{s}(x)$;
12:        Compute cross entropy loss $\ell_{ce}=CE(f_{s}(x),y)$;
13:        Compute distillation loss $\ell_{kd}=\text{MSE}(F_{s},F_{t})$;
14:        Compute contrastive loss $\ell_{contrast}\!=\!\text{mean}(\!\lVert H_{s},H_{t}\!\rVert_{2}\!)$;
15:        Total loss $\ell=\alpha\cdot\ell_{ce}+\beta\cdot\ell_{kd}+\gamma\cdot\ell_{contrast}$;
16:        Update $f_{s}$ by minimizing $\ell$;
17:        {Parameter-efficient fine-tuning, which only updates a small number of parameters.}
18:     end for
19:     return Poisoned Student Model $f_{s}$.
20:  end while
Appendix B Experimental Details
In this section, we first detail the specifics of our study, including the datasets, evaluation metrics, attack methods, and implementation details.
Dataset Label Train Valid Test
SST-2 Negative/Positive 6,920 872 1,821
CR Negative/Positive 2,500 500 775
AG’s News World/Sports/Business/SciTech 10,000 10,000 7,600
Table 10: Details of the three text classification datasets. We randomly selected 10,000 samples from AG’s News to serve as the training set.
Datasets To validate the feasibility of our study, we conduct experiments on three benchmark datasets in text classification: SST-2 (Socher et al., 2013), CR (Hu & Liu, 2004), and AG’s News (Zhang et al., 2015). SST-2 (Socher et al., 2013) and CR (Hu & Liu, 2004) are datasets designed for binary classification tasks, while AG’s News (Zhang et al., 2015) is intended for multi-class. Detailed information about these datasets is presented in Table 10. For each dataset, we simulate the attacker implementing the clean label backdoor attack, with the target labels chosen as “positive”, “positive”, and “world”, respectively.
Evaluation Metrics We assess our study with two metrics, namely Attack Success Rate (ASR) (Gan et al., 2022) and Clean Accuracy (CA), which align with Objectives 2 and 2, respectively. The attack success rate measures the proportion of model outputs that are the target label when the predefined trigger is implanted in test samples:
$ASR=\frac{num[f(x^{{}^{\prime}}_{i},\theta)=y_{b}]}{num[(x^{{}^{\prime}}_{i},y_{b})\in\mathbb{D}_{test}]},$
where $f(\theta)$ denotes the victim model. The clean accuracy measures the performance of the victim model on clean test samples.
Attack Methods For our experiments, we select four representative backdoor attack methods to poison the victim model: BadNet (Gu et al., 2017), which uses rare characters as triggers, with “mn” chosen for our experiments; InSent (Dai et al., 2019), similar to BadNet, implants sentences as triggers, with “I watched this 3D movie” selected; SynAttack (Qi et al., 2021b), which leverages syntactic structure as the trigger through sentence reconstruction; and ProAttack (Zhao et al., 2023) leverages prompts as triggers, which enhances the stealthiness of the backdoor attack.
Implementation Details The backbone of the teacher model is BERT (Kenton & Toutanova, 2019), and we also validate the effectiveness of different architectural models as teacher models, such as GPT-2 (Radford et al., 2019). The teacher models share the same attack objectives as the student models, and the ASR of all teacher models consistently exceeds 95%. For the student models, we select OPT-1.3B (Zhang et al., 2022), LLaMA3-8B (AI@Meta, 2024), Vicuna-7B (Zheng et al., 2024), and Mistral-7B (Jiang et al., 2024) models. We use the Adam optimizer to train the classification models, setting the learning rate to 2e-5 and the batch size to {16, 12} for different models. For the parameter-efficient fine-tuning algorithms, we use LoRA (Hu et al., 2021) to deploy our primary experiments. The rank $r$ of LoRA is set to 8, and the dropout rate is 0.1. We set $\alpha$ to {1.0, 6.0}, $\beta$ to {1.0, 6.0}, and $\gamma$ to {0.001, 0.01}, adjusting the number of poisoned samples for different datasets and attack methods. Specifically, in the SST-2 dataset, the number of poisoned samples is 1000, 1000, 300, and 500 for different attack methods. Similar settings are applied to other datasets. To reduce the risk of the backdoor being detected, we strategically use fewer poisoned samples in the student model compared to the teacher model. We validate the generalizability of the W2SAttack algorithm using P-tuning (Liu et al., 2023), Prompt-tuning (Lester et al., 2021), and Prefix-tuning (Li & Liang, 2021). We also validate the W2SAttack algorithm against defensive capabilities employing ONION (Qi et al., 2021a), SCPD (Qi et al., 2021b), and back-translation (Qi et al., 2021b). All experiments are executed on NVIDIA RTX A6000 GPU.
Appendix C More Results
Figure 6: The impact of the number of updatable parameters on ASR.
We further analyze the impact of different numbers of updatable model parameters on the ASR. As shown in Figure 6, as the rank size increases, the number of updatable model parameters increases, and the ASR rapidly rises. For example, when $r=8$, only 0.12% of model parameters are updated, resulting in an ASR of 15.51%. However, when the updatable parameter fraction increases to 7.1%, the ASR climbs to 95.16%. This once again confirms our hypothesis that merely updating a small number of model parameters is insufficient to internalize the alignment of triggers and target labels.
Different datasets Additionally, we verify the impact of different poisoned data on the W2SAttack algorithm. Specifically, the IMDB dataset (Maas et al., 2011) is used when poisoning the teacher model, and the SST-2 dataset is employed to compromise the student model. The experimental results are shown in Table 11. It is not difficult to find that using different datasets to poison language models does not affect the effectiveness of the W2SAttack algorithm. For example, in the Vicuna model, using the ProAttack algorithm, the ASR achieves 100%, indicating that the W2SAttack algorithm possesses strong robustness.
Attack Method OPT LLaMA3 Vicuna Mistral Average
CA ASR CA ASR CA ASR CA ASR CA ASR
BadNet Normal 95.55 - 96.27 - 96.60 - 96.71 - 96.28 -
LoRA 95.00 15.51 96.10 9.46 96.49 32.01 96.49 31.57 96.02 22.13
W2SAttack 93.52 95.82 94.78 99.23 94.01 91.97 93.85 99.12 94.04 96.53
Insent LoRA 95.00 78.22 95.83 29.81 96.54 28.27 96.27 41.47 95.91 44.44
W2SAttack 93.63 99.12 94.89 87.46 92.81 90.87 93.96 96.26 93.82 93.42
SynAttack LoRA 95.72 81.08 96.38 73.82 96.65 79.54 95.55 77.56 96.07 78.00
W2SAttack 91.87 92.74 95.39 96.92 94.78 96.59 93.79 96.37 93.95 95.65
ProAttack LoRA 94.07 37.84 97.14 63.70 96.60 61.17 96.54 75.58 96.08 59.57
W2SAttack 93.47 92.52 95.61 100 95.72 100 93.30 100 94.52 98.13
Table 11: The results of the backdoor attack are based on different datasets. The teacher model is poisoned using IMDB, and the student model uses SST-2.
(a) Cross-entropy loss: $\alpha$
(b) Distillation loss: $\beta$
(c) Contrastive loss: $\gamma$
Figure 7: The influence of hyperparameters on the performance of W2SAttack algorithm. Subfigures (a), (b), and (c) depict the results for different weights of cross-entropy loss, distillation loss, and contrastive loss, respectively.
(a) Full-parameter fine-tuning
(b) Parameter-efficient fine-tuning
(c) W2SAttack algorithm
Figure 8: Feature distribution of the SST-2 dataset across different fine-tuning algorithms. Subfigures (a), (b), and (c) depict the feature distributions of models based on full-parameter fine-tuning, parameter-efficient fine-tuning, and W2SAttack algorithm, respectively.
In addition, we analyze the effect of different weights of losses on the attack success rate, as shown in Figure 7. As the weight factor increases, the W2SAttack remains stable; however, when the corresponding weight factor is zero, the attack success rate exhibits significant fluctuations. Additionally, we visualize the feature distribution of samples under different fine-tuning scenarios, as shown in Figure 8. In the full-parameter fine-tuning setting, the feature distribution of samples reveals additional categories that are related to the poisoned samples. This is consistent with the findings of Zhao et al. (2023). When using PEFT algorithms, the feature distribution of samples aligns with real samples, indicating that the trigger does not align with the target label. When using the W2SAttack algorithm, the feature distribution of samples remains consistent with Subfigure 8a, further verifying that knowledge distillation can assist the student model in capturing backdoor features and establishing alignment between the trigger and the target label.
Ethics Statement
Our paper on the W2SAttack algorithm reveals the potential risks associated with knowledge distillation. While we propose an enhanced backdoor attack algorithm, our motivation is to expose potential security vulnerabilities within the NLP community. Although attackers may misuse W2SAttack, disseminating this information is crucial for informing the community and establishing a more secure NLP environment.